In a titration process, 32.00 mL of 0.150 M of silver nitrate (AgNO3 (aq)) reacts with 40.00 mL potassium sulfate (K2SO4 (aq) ) . Find the molarity of Potassium Sulfate.

A: 0.905

B: 0.865

C: 0.0743

D: 0.0600
Answer #1

Balanced chemical equation is:
2 AgNO3 + K2SO4 ---> Ag2SO4 + 2 KNO3


Here:
M(AgNO3)=0.15 M
V(AgNO3)=32.0 mL
V(K2SO4)=40.0 mL

According to balanced reaction:
1*number of mol of AgNO3 =2*number of mol of K2SO4
1*M(AgNO3)*V(AgNO3) =2*M(K2SO4)*V(K2SO4)
1*0.15*32.0 = 2*M(K2SO4)*40.0
M(K2SO4) = 0.06 M

Answer: D